OpenStudy (anonymous):

Anyone know how to do this?? For compound interest savings, what are the interest earnings for the lender the first year for $2,300 saved at 13% annual interest compounded annually? A. $199 B. $299 C. $325 D. $330

OpenStudy (tkhunny):

This is a very confusing question. Compounding only one time for exactly one year is exactly the same is Simple Interest. Simple Interest 2300 * 0.13 = 299 Compound 2300 * (1.13) - 2300 = 2599 - 2300 = 299

OpenStudy (anonymous):

Thanks @tkhunny!!